IYF Corporation manufactures miscellaneous parts for building construction and maintenance. IYF uses a normal job costing system. The system applies manufacturing overhead on the basis of direct labor cost. For managerial purposes, over- or underapplied overhead is written off to Cost of Goods Sold monthly. IYF hires interns to work in its Plant Accounting department and, as a part of its interview process, asks candidates to take a short quiz.

You are given the following journal entries for June. (Assume that entries are made only once each month.)

Account Titles Debit Credit
Work-in-Process Inventory (Direct Labor) 10,000
Wages Payable 10,000
Direct Material Inventory 14,700
Accounts Payable 14,700
Finished Goods Inventory 35,700
Work-in-Process Inventory 35,700
Cost of Goods Solda 44,500
Finished Goods Inventory 44,500

aThis entry does not include any over- or underapplied overhead. Over- or underapplied overhead is written off to Cost of Goods Sold once for the month. For June, the amount written off was 4 percent of overhead applied for June.

The Work-in-Process ending account balance on June 30 was twice the beginning balance. The Direct Material Ending Inventory balance on June 30 was $6,930 less than the beginning balance. The Finished Goods ending balance on June 30 was $3,020.

The June income statement shows Cost of Goods Sold of $44,820.

Required:

a. What was the Finished Goods beginning inventory on June 1?

b. How much manufacturing overhead was applied for June?

c. Overhead is applied on the basis of direct labor costs. What was the manufacturing overhead rate for June?

d. How much manufacturing overhead was incurred for June?

e. What was the Work-in-Process beginning inventory balance?

f. What was the Work-in-Process ending inventory balance?
Answer #1
a] Finished Goods beginning Inventory = Finished Goods Sold+ Closing finished goods -Finished Goods completed (WIP transfer)
= 44500+3020-35700 = 11820
b] Overhead Underapplied/(Overapplied ) = Cost of Goods Sold - FG transferred to COGS
= 44820 - 44500 = 320
Manufacturing Overhead applied for June = Underapplied Overhead/ 4%
= 320/4% = 8000
c] Overhead rate based on Direct labour = Manufacturing Overhead applied*100/Direct Labour Cost
=8000*100/10000 = 80% of Direct Labour Cost
d] Manufacturing Overhead incurred = Manufacturing Overhead applied + Underapplied manufacturing overhead
=8000+320 =8320